When sufficient copper and nitric acid reacted to make 645 cm3 of NO (g) @SATP, how many grams of water were produced?
1
Expert's answer
2020-05-08T14:05:30-0400

4HNO3 + 3Cu = 3CuNO3 + NO +2H2O

V=n x 22.4.

n=V/22.4 = 645/22.4 x 1000 = 0,0287946429 moles.

mH2O= 18 x 2 x 0,0287946429= 1,03660714 g of water.